OpenStudy (anonymous):

Prove that: \[cosec^{6}x-\cot^{6}x=3cosec^{2}xcot^{2}x +1\]

OpenStudy (maheshmeghwal9):

\[(cosec^2 x)^3-(\cot^2x)^3.\] Use \[a^3-b^3=(a-b){(a^2+ab+b^2)}.\] http://in.answers.yahoo.com/question/index?qid=20101115204133AAm7phA

OpenStudy (maheshmeghwal9):

the link would provide help:D

OpenStudy (anonymous):

on that link,the ques. involves 1 + 3 cot^2x+3cot^4x and check my ques.

OpenStudy (anonymous):

@mukushla

OpenStudy (anonymous):

1 + 3 cot^2x+3cot^4x=1+3cot^2x(1+cot^2x)=1+3(cot^2x)(cosec^2x)

OpenStudy (maheshmeghwal9):

@shubham.bagrecha I didn't provided that link for whole solution. I only wanted to help so that u can prove ur question further as @mukushla did :)